# Designing lattice spin models and magnon gaps with supercurrents

### v1

@@ III.1 Equal spin pairing
where $A,B$ denote the sublattices of the antiferromagnet. We assume for simplicity that the $s-d$-coupling strength $\Lambda_{l}=\Lambda$ is the same for both sublattices. For a superconductor with equal-spin Cooper pairs, the coherence factors $u_{k,\sigma}$ and $v_{k,\sigma}$ that diagonalize the Hamiltonian are given in Eq. (S4). The corresponding eigenvalues $E_{k,\sigma}$ are given in Eq. (S5). Rewriting the $c$–operators in terms of the $\gamma$–operators according to Eq. (S3) and taking the average over the $\gamma$–operators gives
@@
If the triplet superconducting OP $|\Delta_{k}^{\sigma}|=|\Delta_{k}^{-\sigma}|$, the magnon gap correction $\mathcal{D}$ is zero in the presence of a pure charge supercurrent or a pure spin supercurrent. For the charge current $Q^{\sigma}=Q$, the absolute value of the coherence factors in Eq. (S4) and the eigenenergies in Eq. (S5) are spin independent. Therefore, the spin sum in Eq. (S18) is zero. For a pure spin supercurrent $Q^{\sigma}=-Q^{-\sigma}$, we use that $\epsilon_{k}=\epsilon_{-k}$. Again, we find that the absolute value of the coherence factors in Eq. (S4) are spin independent, and the eigenenergy $E_{k,\sigma}=E_{-k,-\sigma.}$. Since $|u_{k}|=|u_{-k}|$ and $|v_{k}|=|v_{-k}|$, we find again that the spin sum and thus $\mathcal{D}$ is zero. A non-zero $\mathcal{D}$ is thus obtained for a spin-polarized charge supercurrent ($Q^{\uparrow}\neq\pm Q^{\downarrow}$).
III.2 Spin singlet $s$-wave Cooper pairs
The Hubbard Hamiltonian in real space for a spin-split spin singlet $s$-wave superconductor is
@@ III.2 Spin singlet $s$-wave Cooper pairs
This gives, as before, magnon energies $\omega_{\alpha,k}=\omega^{\prime}_{\alpha,k}+\mathcal{D}^{s}$ $\omega_{\beta,k}=\omega^{\prime}_{\beta,k}-\mathcal{D}^{s}$, where $\omega^{\prime}_{\alpha,k}$ and $\omega^{\prime}_{\beta,k}$ do not depend on the supercurrent. The magnon gap correction is
@@
The magnitudes of the coherence factors are spin independent and the eigenenergy depends on spin only through the spin splitting field. Therefore, a non-zero spin splitting is required to achieve a nonzero magnon gap correction. The magnon gap correction can be simplified by shifting $\sum_{k}\rightarrow\sum_{k+Q}$ and relabeling $k,\sigma\rightarrow-k,-\sigma$ in the sum over the $v$–term:
@@
IV Gap equation for the superconducting order parameter
In the numerical calculation of the coupling constants in Eq. (S13), we assumed that the magnitude of the superconducting OP was independent of the applied supercurrent. This is justified in this section. Moreover, when we calculated the supercurrent-induced contribution to the magnon gap, we assumed that the magnitude of the superconducting OP at $T=0$, $h=0$ and in the absence of any supercurrents was $\Delta_{0}^{\sigma}=0.1t$. At finite temperature, the magnitude of the OP is reduced and depends on the supercurrent, and it is calculated by solving the gap equation.
IV.1 Equal spin pairing
@@
Figure S4(a) shows the $Q^{\sigma}$-dependence of the OP magnitude. Below the critical current, it is reasonable to approximate the gap as $Q^{\sigma}$-independent. Therefore, it is reasonable to approximate the OP magnitude as spin independent since the OP magnitude depends on spin only through the Cooper pair momentum $Q^{\sigma}$. The critical Cooper pair momentum is $aQ_{c}\approx 0.05\approx 2\pi\cdot 8/1000$ at $T=0$ and $aQ_{c}\approx 0.035\approx 2\pi\cdot 6/1000$ at $T=0.2T_{c}$. Here, $T_{c}\approx 0.05t/k_{B}$ is the critical temperature when $Q^{\sigma}=0$, as shown in Fig. S4(b). The supercurrent magnitude $aQ\approx 0.044$ used in Fig. 1 and 2 in the main text is below the critical supercurrent at $T=0$.
@@
As for realistic materials, a critical current $aQ_{c}=0.05$ is quite large. Take for example NbTi, which is a type II superconductor known to have a high critical current. The critical current is given by $Q_{c}=\frac{2\Delta_{0}}{\hbar v_{F}}$, and the BCS coherence length is $\xi_{0}=\frac{\hbar v_{F}}{\pi\Delta_{0}}.$ This gives $Q_{c}={2}/{(\pi\xi_{0})}$ Tinkham (2004). The coherence length of NbTi is $\xi_{0}/a\approx 0.2$, which gives $aQ_{c}\approx 0.04$. Even though the superconducting OP magnitude and critical current we use are large, we argue that it is reasonable to use these values. The effects predicted by second-order perturbation theory are expected to be small, and using a large OP and critical current is a way of allowing the effects to be larger even within the perturbative approach.
IV.2 Spin singlet $s$-wave Cooper pairs
@@
The gap equation is solved numerically in Fig. S5. The supercurrent magnitudes used in Fig. 3(d) in the main text are well below the critical current, as shown in Fig. S5(a). The temperatures are normalized on the critical temperature $T_{c}\approx 0.057t/k_{B}$ at $Q=0$ and $h=0$, as shown in Fig. S5(b).

### v2

@@ IV.1 Equal spin pairing
where $A,B$ denote the sublattices of the antiferromagnet. We assume for simplicity that the sds-d-coupling strength $\Lambda_{l}=\Lambda$ is the same for both sublattices. For a superconductor with equal-spin Cooper pairs, the coherence factors $u_{k,\sigma}$ and $v_{k,\sigma}$ that diagonalize the Hamiltonian are given in Eq. (S4). The corresponding eigenvalues $E_{k,\sigma}$ are given in Eq. (S5). Rewriting the cc–operators in terms of the $\gamma$–operators according to Eq. (S3) and evaluating the ground-state quantum expectation value for the $\gamma$–operators gives
@@
If the triplet superconducting OP has equal magnitude for both spins, $|\Delta_{k}^{\sigma}|=|\Delta_{k}^{-\sigma}|$, the magnon gap correction $\mathcal{D}$ is zero in the presence of a pure charge supercurrent or a pure spin supercurrent. For the charge current $Q^{\sigma}=Q$, the absolute value of the coherence factors in Eq. (S4) and the eigenenergies in Eq. (S5) are spin independent. Therefore, the spin sum in Eq. (S22) is zero. For a pure spin supercurrent $Q^{\sigma}=-Q^{-\sigma}$, we use that $\epsilon_{k}=\epsilon_{-k}$. Again, we find that the absolute value of the coherence factors in Eq. (S4) are spin independent, and the eigenenergy $E_{k,\sigma}=E_{-k,-\sigma.}$. Since $|u_{k}|=|u_{-k}|$ and $|v_{k}|=|v_{-k}|$, we find again that the spin sum and thus $\mathcal{D}$ is zero. A non-zero $\mathcal{D}$ is thus obtained for a spin-polarized charge supercurrent ($Q^{\uparrow}\neq\pm Q^{\downarrow}$).
IV.2 Spin singlet $s$-wave Cooper pairs
The attractive Hubbard Hamiltonian in real space for a spin-split spin singlet $s$-wave superconductor is
@@ IV.2 Spin singlet $s$-wave Cooper pairs
This gives, as before, magnon energies $\omega_{\alpha,k}=\omega^{\prime}_{\alpha,k}+\mathcal{D}^{s}$ and $\omega_{\beta,k}=\omega^{\prime}_{\beta,k}-\mathcal{D}^{s}$, where $\omega^{\prime}_{\alpha,k}$ and $\omega^{\prime}_{\beta,k}$ do not depend on the supercurrent. The magnon gap correction is
@@
The magnitudes of the coherence factors are spin-independent and the eigenenergy depends on spin only through the spin splitting field. Therefore, a non-zero spin splitting is required to achieve a nonzero magnon gap correction. The magnon gap correction can be simplified by shifting $\sum_{k}\rightarrow\sum_{k+Q}$ and relabeling $k,\sigma\rightarrow-k,-\sigma$ in the sum over the vv–term:
@@
V Gap equation for the superconducting order parameter
In this section, we identify the critical supercurrent magnitude $Q_{c}$ from solving the gap equation. This value is used to normalize the supercurrent strengths in the figures in the main text and in the Supplemental material. Moreover, we solve the gap equation to determine how the magnitude $\Delta^{\sigma}$ of the superconducting order parameter depends on the applied supercurrent, the spin splitting and the temperature. The resulting value is used in the calculations for the coupling constants in Eq. (S13) and in the magnon gap corrections in Eqs. (S22) and (S39).
We underline that we consider an effective 1D model for simplicity since this allows us to numerically simulate larger systems, whereas the actual experimental realization of our setup would be in 3D (a thin-film geometry). It is well-known that effective 1D models can describe experimental systems in 3D well when physical quantities primarily vary along one coordinate axis le Sueur et al. (2008).
V.1 Equal spin pairing
@@
Figure S6(a) shows the $Q^{\sigma}$-dependence of the OP magnitude. Below the critical current, it is reasonable to approximate the gap as $Q^{\sigma}$-independent. Therefore, it is reasonable to approximate the OP magnitude as spin independent since the OP magnitude depends on spin only through the Cooper pair momentum $Q^{\sigma}$. The critical Cooper pair momentum is $aQ_{c}\approx 0.0505\approx 2\pi\cdot 8/1000$ at $T=0$ and $aQ_{c}\approx 0.035\approx 2\pi\cdot 6/1000$ at $T=0.2T_{c}$. Here, $T_{c}\approx 0.05t/k_{B}$ is the critical temperature when $Q^{\sigma}=0$, as shown in Fig. S6(b). The supercurrent magnitude $aQ\approx 0.044$ used in Fig. 1 and 2 in the main text is below the critical supercurrent at $T=0$.
@@
V.2 Spin singlet $s$-wave Cooper pairs
@@
The gap equation is solved numerically in Fig. S7. The supercurrent magnitudes used in Fig. 3(d) in the main text are normalized on the critical current $Q_{c}$ at which the superconducting order parameter drops to zero, see Fig. S7(a). The temperatures are normalized on the critical temperature $T_{c}\approx 0.057t/k_{B}$ at $Q=0$ and $h=0$, as shown in Fig. S7(b).
VI Tuning the supercurrents
The momentum $Q$ of the Cooper pairs is related to the supercurrent $j$ through $j=-2ten_{e}a^{2}Q/\hbar$ for $|Qa|\ll 1$ and at low temperature Takashima et al. (2017). Here, $n_{e}$ is the density of superconducting electrons. While there are several setups that will allow for charge and spin current injection into a superconductor, which are both converted into a supercurrent when triplet pairs are present, we provide one such example in Fig. S8. In (a), the impurity spins are placed on the surface of the superconductor and their ground state configuration is controlled by the supercurrent. In (b), an antiferromagnetic or altermagnetic insulating film is grown on top of the superconductor and the supercurrent controls the magnon gap. The superconductor is connected to two normal electrodes (gray), and the voltage $V_{1}$ controls the magnitude of the injected charge current $j_{c}$. This resistive current is converted to a supercurrent by the order parameter over the charge relaxation length $\lambda_{Q}$. The spin supercurrent is injected from the Hall bar (blue) by injection of a spin accumulation into the superconductor where the triplet order parameter converts it to a spin supercurrent. The voltage $V_{2}$ controls the magnitude of the transverse charge current and thus the spin supercurrent magnitude. As long as the contact between the Hall bar and the superconductor is smaller than $\lambda_{Q}$, the current will not be converted to a supercurrent in the superconductor and the Hall bar will therefore not be short-circuited. Moreover, the charge supercurrent $j_{c}$ will not escape into the Hall bar. If $V_{1}\neq 0$ and $V_{2}=0$, a pure charge supercurrent flows in the superconductor. If $V_{1}=0$ and $V_{2}\neq 0$, a pure spin supercurrent flows in the superconductor.
Figure S8: Proposed setup for controlling the supercurrent in the superconductor. In (a), the ground state configuration of two magnetic adatoms are controlled by the supercurrent. In (b), the magnon gap in a thin-film antiferromagnetic or altermagnetic insulator (AFMI) is controlled by the supercurrent.